merchandise exports= 82.1, merchandise imports= 128.7, invisibles= 29.9, services= 18.7, transfers= 17.0, income[net]= -5.8, capital accounts= 4.8, errors= -0.3. calculate trade balance


1
Expert's answer
2022-03-27T18:55:03-0400

Trade balance is the difference between the country's exports and imports.

therefore;

"BOP = X-M"


where "X" - Exports

"M" - Imports

"BOP = 82.1-128.7 = -46.6"

The BOP is negative and therefore there is a trade deficit.